Bromine has two isotopes, Br 79 and Br 81. The isotopes occur in a 50:50 ratio. Given that the mass spectrum of bromine contains peaks for both bromine atoms and diatomic bromine molecules, predict the number of peaks in the spectrum. What would be the relative height of the atomic peaks? What would be the relative height of the molecular peaks? Assume that z = +1 in each case.
1
Expert's answer
2019-08-20T08:19:26-0400

Relative height of atomic peaks m/z= 79Br-79Br+ =158, 79Br-81Br+ =160 and 81Br-81Br+ = 162

Relative height of the molecular peaks= (50 "\times" 79)+(50"\times" 81)= 8000= "\frac{8000}{100}" = 80